# DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators

### v1

@@ Abstract
Generally, the decoder-only large language models (LLMs) are adapted to context-aware neural machine translation (NMT) in a concatenating way, where LLMs take the concatenation of the source sentence (i.e., intra-sentence context) and the inter-sentence context as the input, and then to generate the target tokens sequentially. This adaptation strategy, i.e., concatenation mode, considers intra-sentence and inter-sentence contexts with the same priority, despite an apparent difference between the two kinds of contexts. In this paper, we propose an alternative adaptation approach, named Decoding-enhanced Multi-phase Prompt Tuning (DeMPT), to make LLMs discriminately model and utilize the inter- and intra-sentence context and more effectively adapt LLMs to context-aware NMT. First, DeMPT divides the context-aware NMT process into three separate phases. During each phase, different continuous prompts are introduced to make LLMs discriminately model various information. Second, DeMPT employs a heuristic way to further discriminately enhance the utilization of the source-side inter- and intra-sentence information at the final decoding phase. Experiments show that our approach significantly outperforms the concatenation method, and further improves the performance of LLMs in discourse modeling.111As this paper is under review, we will release our code and datasets later.
@@ 1 Introduction
Context-aware neural machine translation (NMT) goes beyond sentence-level NMT by incorporating inter-sentence context at the document level Zhang et al. (2018); Miculicich et al. (2018); Voita et al. (2018, 2019b, 2019a); Bao et al. (2021); Sun et al. (2022), aiming to address discourse-related challenges such as zero pronoun translation Wang et al. (2019), lexical translation consistency Lyu et al. (2021, 2022), and discourse structure Hu and Wan (2023). A recent paradigm shift has been witnessed in context-aware NMT with the emergence of the decoder-only large language models (LLMs) BigScience (2022); Google (2022); MetaAI (2023b, a); OpenAI (2023). These generative language models, trained on extensive public data, have gained significant attention in the natural language processing (NLP) community. In adapting LLMs to context-aware NMT, a common strategy involves concatenating multiple source sentences as a prefix and generating translations token-by-token, relying on the prefix and previously predicted target tokens, as shown in Figure 1 (a). However, a critical observation of this strategy reveals a potential drawback – the equal prioritization of the inter- and intra-sentence contexts during token generation. Importantly, the intra-sentence context inherently contains richer parallel semantic information with the target sentence and should be given a higher priority than the inter-sentence context. Consequently, we propose that separately modeling and utilizing the inter- and intra-sentence contexts should explicitly inform LLMs of the document-level context and the current sentence itself, thus being able to prevent the misallocation of attention weights to source-side tokens Bao et al. (2021); Li et al. (2023). Inspired by the success of prompt tuning Li and Liang (2021); Liu et al. (2022); Tan et al. (2022), our alternative approach, named Decoding-Enhanced Multi-phase Prompt Tuning (DeMPT), aims to enhance LLMs’ adaptability to context-aware NMT, as shown in Figure 1 (b).
Specifically, we divide the whole procedure of context-aware NMT into three phases: inter-sentence context encoding, intra-sentence context encoding, and decoding. Following Li and Liang (2021); Liu et al. (2022), we sequentially and differentially adapt LLMs for each phase, utilizing phase-specific trainable prompts. This phased tuning method enables LLMs to independently capture and model both inter- and intra-sentence contexts, facilitating a better understanding of their differences. Importantly, our approach only divides the original input into three parts without significantly increasing computational load. As a result, there is no substantial decrease in inference speed compared to the concatenating method, as detailed in Section 4.3.
@@
In summary, our contributions can be outlined as follows:
We propose a novel multi-phase prompt tuning approach to divide context-aware NMT into three phases, making LLMs aware of the distinction between inter- and intra-sentence contexts.
@@
We validate our approach using llama-2-7b and bloomz-7b1-mt as foundation models, demonstrating its effectiveness across five context-aware translation directions. Extensive analyses further highlight the substantial enhancement in LLMs’ ability for context-aware NMT.
@@ 2.1 Multi-phase Encoding and Decoding
where is the output of the -th transformer block. FFN and Multi-Attn are the feed-forward network sublayer and multi-head self-attention sublayer, respectively.222For simplicity, we omit the normalization and residual operations in this paper. and are the concatenating and slicing operations, respectively.
@@ 2.2 Enhanced Decoding Phase
Therefore, to address above two issues, we propose an enhanced decoding phase with an aim to more effectively utilize both the inter- and intra-sentence contexts. Inspired by Kuang et al. (2018), we move both the two types of inter- and intra-sentence contexts closer to target words to achieve a tight interaction between them. Specifically, we concatenate the decoding states with the two types of representations to predict the next target words. As shown in Figure 3, the enhanced next word prediction is a combination of three distributions with different inputs:
where and control the contribution of and , respectively, which can be further computed as:
where is same as in Eq. 16, and are the averaged and at token level, respectively.
@@ 2.3 Phase-aware Prompts
We emphasize the LLM needs to play various roles across three phases, and maintaining similar prompts across different phases may not be reasonable. Thus, we empower LLM to distinguish different phases by introducing a type embedding and a transfer layer333Different from the multi-layer perceptron (MLPs) used for reparameterization, our transfer layer is shared-parameter for all prompts. Thus, there are fewer trainable parameters during the training of our model. We compare the number of trainable parameters among different tuning methods in Table 3 and analyze the effect of the transfer layer in Appendix D. for these prompts:
where is randomly initialized prompt, are trainable parameters, and TypeEmb() is type embeddings layer of prompts. represents either phase 1, phase 2, or phase 3.
@@ 3 Experimentation
| Model | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | |
| Trans. | 29.86 | 0.8406 | 38.53 | 0.8545 | 41.44 | 0.8682 | 48.74 | 0.8783 | 32.25 | 0.8169 |
| llama-2-7b as foundation model | ||||||||||
| MT-LoRA | 27.43 | 0.8511 | 38.18 | 0.8647 | 40.96 | 0.8712 | 47.52 | 0.8733 | 33.00 | 0.8311 |
| MT-PT | 31.32 | 0.8565 | 41.92 | 0.8675 | 43.56 | 0.8752 | 51.32 | 0.8819 | 35.46 | 0.8333 |
| CMT-PT | 31.13 | 0.8387 | 42.01 | 0.8699 | 43.11 | 0.8762 | 51.66 | 0.8823 | 35.91 | 0.8396 |
| MPT | *33.21 | 0.8645 | †43.11 | 0.8744 | *43.88 | 0.8824 | †52.01 | 0.8913 | †36.49 | 0.8456 |
| DeMPT | *33.89 | 0.8658 | †43.71 | 0.8816 | *44.69 | 0.8899 | †53.10 | 0.8979 | †36.55 | 0.8438 |
| bloomz-7b1-mt as foundation model | ||||||||||
| MT-LoRA | 25.79 | 0.8466 | 35.67 | 0.8601 | 35.17 | 0.8522 | 46.32 | 0.8644 | 28.01 | 0.8012 |
| MT-PT | 30.99 | 0.8520 | 40.49 | 0.8661 | 37.76 | 0.8579 | 50.68 | 0.8823 | 30.27 | 0.8106 |
| CMT-PT | 30.82 | 0.8504 | 40.31 | 0.8639 | 38.01 | 0.8601 | 50.26 | 0.8832 | 29.80 | 0.8108 |
| MPT | *31.81 | 0.8601 | *41.11 | 0.8766 | †38.99 | 0.8669 | *51.33 | 0.8910 | *30.99 | 0.8201 |
| DeMPT | *32.46 | 0.8649 | *41.92 | 0.8790 | †40.06 | 0.8703 | *52.25 | 0.8990 | *31.79 | 0.8253 |
| Model | ZH | FR | DE | ES | RU |
| Trans. | 47.63 | 54.41 | 58.29 | 62.52 | 48.79 |
| llama-2-7b as foundation model | |||||
| MT-LoRA | 44.83 | 54.52 | 57.72 | 62.18 | 49.06 |
| MT-PT | 49.49 | 57.87 | 60.89 | 65.02 | 52.59 |
| CMT-PT | 49.53 | 58.27 | 61.23 | 65.89 | 53.34 |
| MPT | 51.56 | 59.56 | 62.15 | 67.14 | 54.18 |
| DeMPT | 52.68 | 60.33 | 63.11 | 67.95 | 54.34 |
| bloomz-7b1-mt as foundation model | |||||
| MT-LoRA | 43.23 | 51.82 | 51.12 | 61.77 | 43.29 |
| MT-PT | 49.48 | 56.81 | 55.40 | 64.71 | 46.14 |
| CMT-PT | 49.61 | 57.05 | 55.81 | 65.12 | 46.09 |
| MPT | 50.22 | 57.93 | 56.69 | 66.25 | 47.29 |
| DeMPT | 50.62 | 58.30 | 57.34 | 67.12 | 48.00 |
We build our approach upon two open-source LLMs, namely, llama-2-7b444https://huggingface.co/meta-llama/Llama-2-7b-hf and bloomz-7b1-mt555https://huggingface.co/bigscience/bloomz-7b1-mt. We verify the effectiveness of our proposed approach on five translation tasks, including {Chinese (ZH), French (FR), German (DE), Spanish (ES), Russian (RU)}English (EN).
@@ 3.1 Experimental Settings
The corpus of all translation tasks is extracted from New-Comentary-v18. See Appendix A for splitting and statistics of the training set, valid set, and test set. We use the tokenizer of foundation models to process the input data and no any other preprocessing is performed.
@@
We compare our approach against four baselines:
Transformer (Trans.): It is an encoder-decoder Transformer-base model Vaswani et al. (2017) that is trained from scratch.
MT-LoRA: It is a tuned LLM adapted to NMT task via the tuning method of Low-Rank Adaptation Hu et al. (2022), which makes large-scale pre-training models adapt to a new task by injecting a trainable rank decomposition matrice into each layer of the Transformer architecture.
MT-PT: It is a tuned LLM adapted to NMT task via the deep prompt tuning with MLPs reparameterization,666We attempt to remove reparameterization but experience a significant decline in performance. which only tunes continuous prompts with a frozen language model.
CMT-PT: Similar to MT-PT, it is also a tuned LLM via the deep prompt tuning with MLPs reparameterization. Unlike MT-PT, it utilizes inter-sentence context within the concatenation strategy, as depicted in Figure 1 (a).
Among them, Transformer, MT-LoRA, and MT-PT are context-agnostic systems while CMT-PT is a context-aware system. For a fair comparison, we ensure that all context-aware systems, including CMT-PT, MPT, and DeMPT, incorporate identical inter-sentence context.
@@
For the Transformer model, we implement it upon Fairseq Ott et al. (2019). For MT-LoRA models, we set the rank of trainable matrices as 16 which performs best in our preliminary experiment. For all MT-PT models, CMT-PT models, and our models, we set the prompt length as 64. For the incorporation of inter-sentence context in CMT-PT models and our models, we consider a dynamic , in which the total tokens are no more than 256. In enhanced decoding, we consider the three next word predictions to be equally important by setting both and to 1/3. More details of training are provided in Appendix B.
@@
We use sacreBLEU (accuracy-related metric)777Signature: nrefs:1|case:mixed|eff:no|tok:13a|
smooth:exp|version:2.3.1 Post (2018), COMET (semantics-related metric) with the wmt22-comet-da model888https://github.com/Unbabel/COMET Rei et al. (2020), and BlonDe (discourse-related metric) Jiang et al. (2022) as the evaluation metrics.

### v2

@@ Abstract
Generally, the decoder-only large language models (LLMs) are adapted to context-aware neural machine translation (NMT) in a concatenating way, where LLMs take the concatenation of the source sentence (i.e., intra-sentence context) and the inter-sentence context as the input, and then to generate the target tokens sequentially. This adaptation strategy, i.e., concatenation mode, considers intra-sentence and inter-sentence contexts with the same priority, despite an apparent difference between the two kinds of contexts. In this paper, we propose an alternative adaptation approach, named Decoding-enhanced Multi-phase Prompt Tuning (DeMPT), to make LLMs discriminately model and utilize the inter- and intra-sentence context and more effectively adapt LLMs to context-aware NMT. First, DeMPT divides the context-aware NMT process into three separate phases. During each phase, different continuous prompts are introduced to make LLMs discriminately model various information. Second, DeMPT employs a heuristic way to further discriminately enhance the utilization of the source-side inter- and intra-sentence information at the final decoding phase. Experiments show that our approach significantly outperforms the concatenation method, and further improves the performance of LLMs in discourse modeling.111We release our code and datasets on https://github.com/xllyu-nlp/DeMPT.
DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators
Xinglin Lyu♣, Junhui Li♣††thanks: Corresponding author: Junhui Li., Yanqing Zhao♠, Min Zhang♠, Daimeng Wei♠, Shimin Tao♠, Hao Yang♠, Min Zhang♣ ♣School of Computer Science and Technology, Soochow University, Suzhou, China ♠Huawei Translation Services Center, Beijing, China xllv2020@stu.suda.edu.cn, {lijunhui,minzhang}@suda.edu.cn {zhaoyanqing,zhangmin186,weidaimeng,taoshimin,yanghao30}@huawei.com
@@ 1 Introduction
Context-aware neural machine translation (NMT) goes beyond sentence-level NMT by incorporating inter-sentence context at the document level Zhang et al. (2018); Miculicich et al. (2018); Voita et al. (2018, 2019b, 2019a); Bao et al. (2021); Sun et al. (2022), aiming to address discourse-related challenges such as zero pronoun translation Wang et al. (2019), lexical translation consistency Lyu et al. (2021, 2022), and discourse structure Hu and Wan (2023). A recent paradigm shift has been witnessed in context-aware NMT with the emergence of the decoder-only large language models (LLMs) BigScience (2022); Google (2022); MetaAI (2023b, a); OpenAI (2023). These generative language models, trained on massive data, have gained significant attention in the natural language processing (NLP) community. In adapting LLMs to context-aware NMT, a common strategy involves concatenating multiple source sentences as a prefix and generating translations token-by-token, relying on the prefix and previously predicted target tokens, as shown in Figure 1 (a). However, a critical observation of this strategy reveals a potential drawback – the equal prioritization of the inter- and intra-sentence contexts during token generation. Importantly, the intra-sentence context inherently contains richer parallel semantic information with the target sentence and should be given a higher priority than the inter-sentence context. Consequently, we propose that separately modeling and utilizing the inter- and intra-sentence contexts should explicitly inform LLMs of the document-level context and the current sentence itself, thus being able to prevent the misallocation of attention weights to source-side tokens Bao et al. (2021); Li et al. (2023). Inspired by the success of prompt tuning Li and Liang (2021); Liu et al. (2022); Tan et al. (2022), our alternative approach, named Decoding-Enhanced Multi-phase Prompt Tuning (DeMPT), aims to enhance LLMs’ adaptability to context-aware NMT, as shown in Figure 1 (b).222Following the findings of Bao et al. (2021), which indicate that source-side context is relatively more important for document-level MT compared to target-side context, we focus exclusively on source-side context in this paper. Nonetheless, we provide an additional discussion on integrating target-side context in Appendix M.
Specifically, we divide the whole procedure of context-aware NMT into three phases: inter-sentence context encoding, intra-sentence context encoding, and decoding. Following Li and Liang (2021); Liu et al. (2022), we sequentially and differentially adapt LLMs for each phase, utilizing phase-specific trainable prompts. This phased tuning method enables LLMs to independently capture and model both inter- and intra-sentence contexts, facilitating a better understanding of their differences. Our approach splits the input into three parts without significantly increasing computational load, thus maintaining inference speed comparable to concatenation, as detailed in Appendix D.
@@
Our contributions can be summarized as follows:
We introduce a multi-phase prompt tuning approach that divides context-aware NMT into three phases, enabling LLMs to distinguish between inter- and intra-sentence contexts.
@@
We validate our approach using llama-2-7b and bloomz-7b1-mt as foundation models, demonstrating its effectiveness across five translation directions. Extensive analyses further highlight the substantial enhancement in LLMs’ ability for context-aware MT.
@@ 2.1 Multi-phase Encoding and Decoding
where is the output of the -th transformer block. FFN and Multi-Attn are the feed-forward network and multi-head self-attention sublayers, respectively.333For simplicity, we omit the normalization and residual operations in this paper. and are the concatenating and slicing operations, respectively.
@@ 2.2 Enhanced Decoding Phase
Therefore, to address above two issues, we propose an enhanced decoding phase with an aim to more effectively utilize both the inter- and intra-sentence contexts. Inspired by Kuang et al. (2018), we move both the two types of inter- and intra-sentence contexts closer to target words to achieve a tight interaction between them. Specifically, we concatenate the decoding states with the two types of representations to predict the next target words. As shown in Figure 3, the enhanced next word prediction is a combination of three distributions:
where and control the contribution of and , respectively, which can be further formulated as:
where is same as in Eq. 16, and are the averaged and at token level, respectively.444Notably, the computation of and does not require a full decoding forward pass. It involves solely an FFN layer (two linear transformation layers and a ReLU activation layer), an LLM-Head layer (a linear transformation layer), and a softmax function layer. To further identify the effect of inter- and intra-sentence context in this strategy, we provide an ablation study about and in Appendix J.
@@ 2.3 Phase-aware Prompts
We emphasize the LLM needs to play various roles across three phases, and maintaining similar prompts across different phases may not be reasonable. Thus, we empower LLM to distinguish different phases by introducing a type embedding and a transfer layer555Unlike the multi-layer perceptrons (MLPs) used for reparameterization, our transfer layer shares parameters across all prompts, reducing the number of trainable parameters. Table 3 compares the trainable parameters of various tuning methods, and Appendix J analyzes the effect of the transfer layer. for these prompts:
where is randomly initialized prompt, are the trainable parameters, and TypeEmb() is type embeddings layer of the prompts. represents either phase 1, phase 2, or phase 3.
@@ 3 Experimentation
| Model | ZHEN | FREN | DEEN | ESEN | RUEN | Average | ||||||
| BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | BLEU | COMET | |
| ⊘Trans. | 29.86 | 0.8406 | 38.53 | 0.8545 | 41.44 | 0.8682 | 48.74 | 0.8783 | 32.25 | 0.8169 | 38.16 | 0.8517 |
| Traditional context-aware NMT models | ||||||||||||
| ⊙MR-Trans. | 30.61 | 0.8413 | 38.72 | 0.8533 | 42.11 | 0.8693 | 49.69 | 0.8812 | 33.27 | 0.8211 | 38.88 | 0.8532 |
| + mBART | 32.69 | 0.8601 | 42.01 | 0.8759 | 44.61 | 0.8840 | 51.67 | 0.8831 | 36.39 | 0.8459 | 41.39 | 0.8698 |
| ⊙G-Trans. | 30.99 | 0.8411 | 38.96 | 0.8524 | 42.46 | 0.8658 | 49.68 | 0.8794 | 33.59 | 0.8201 | 39.14 | 0.8518 |
| + mBART | 32.99 | 0.8597 | 42.02 | 0.8764 | 44.81 | 0.8836 | 52.07 | 0.8911 | 36.83 | 0.8461 | 41.74 | 0.8714 |
| llama-2-7b as foundation model | ||||||||||||
| ⊘MT-LoRA | 27.43 | 0.8511 | 38.18 | 0.8647 | 40.96 | 0.8712 | 47.52 | 0.8733 | 33.00 | 0.8311 | 37.42 | 0.8583 |
| ⊘MT-PT | 31.32 | 0.8565 | 41.92 | 0.8675 | 43.56 | 0.8752 | 51.32 | 0.8819 | 35.46 | 0.8333 | 40.72 | 0.8629 |
| ⊙CMT-PT | 31.13 | 0.8387 | 42.01 | 0.8699 | 43.11 | 0.8762 | 51.66 | 0.8823 | 35.91 | 0.8396 | 40.76 | 0.8613 |
| ⊙MPT | *33.21 | 0.8645 | †43.11 | 0.8744 | *43.88 | 0.8824 | †52.01 | 0.8913 | †36.49 | 0.8456 | 41.74 | 0.8716 |
| ⊙DeMPT | *33.89 | 0.8658 | †43.71 | 0.8816 | *44.69 | 0.8899 | †53.10 | 0.8979 | †36.55 | 0.8438 | 42.39 | 0.8758 |
| bloomz-7b1-mt as foundation model | ||||||||||||
| ⊘MT-LoRA | 25.79 | 0.8466 | 35.67 | 0.8601 | 35.17 | 0.8522 | 46.32 | 0.8644 | 28.01 | 0.8012 | 34.21 | 0.8449 |
| ⊘MT-PT | 30.99 | 0.8520 | 40.49 | 0.8661 | 37.76 | 0.8579 | 50.68 | 0.8823 | 30.27 | 0.8106 | 38.04 | 0.8539 |
| ⊙CMT-PT | 30.82 | 0.8504 | 40.31 | 0.8639 | 38.01 | 0.8601 | 50.26 | 0.8832 | 29.80 | 0.8108 | 37.84 | 0.8537 |
| ⊙MPT | *31.81 | 0.8601 | *41.11 | 0.8766 | †38.99 | 0.8669 | *51.33 | 0.8910 | *30.99 | 0.8201 | 38.85 | 0.8629 |
| ⊙DeMPT | *32.46 | 0.8649 | *41.92 | 0.8790 | †40.06 | 0.8703 | *52.25 | 0.8990 | *31.79 | 0.8253 | 39.70 | 0.8677 |
| Model | ZH | FR | DE | ES | RU | Avg. |
| ⊘Trans. | 47.63 | 54.41 | 58.29 | 62.52 | 48.79 | 54.33 |
| Traditional context-aware NMT model | ||||||
| ⊙MR-Trans. | 48.51 | 55.55 | 59.02 | 63.51 | 49.88 | 55.29 |
| + mBART | 50.66 | 58.01 | 61.99 | 66.01 | 54.11 | 58.15 |
| ⊙G-Trans. | 48.99 | 55.31 | 59.23 | 63.99 | 50.09 | 55.52 |
| + mBART | 50.98 | 57.88 | 61.97 | 66.21 | 54.33 | 58.27 |
| llama-2-7b as foundation model | ||||||
| ⊘MT-LoRA | 44.83 | 54.52 | 57.72 | 62.18 | 49.06 | 53.66 |
| ⊘MT-PT | 49.49 | 57.87 | 60.89 | 65.02 | 52.59 | 57.17 |
| ⊙CMT-PT | 49.53 | 58.27 | 61.23 | 65.89 | 53.34 | 57.65 |
| ⊙MPT | 51.56 | 59.56 | 62.15 | 67.14 | 54.18 | 58.92 |
| ⊙DeMPT | 52.68 | 60.33 | 63.11 | 67.95 | 54.94 | 59.80 |
| bloomz-7b1-mt as foundation model | ||||||
| ⊘MT-LoRA | 43.23 | 51.82 | 51.12 | 61.77 | 43.29 | 50.25 |
| ⊘MT-PT | 49.48 | 56.81 | 55.40 | 64.71 | 46.14 | 54.51 |
| ⊙CMT-PT | 49.61 | 57.05 | 55.81 | 65.12 | 46.09 | 54.74 |
| ⊙MPT | 50.22 | 57.93 | 56.69 | 66.25 | 47.29 | 55.68 |
| ⊙DeMPT | 50.62 | 58.30 | 57.34 | 67.12 | 48.00 | 56.28 |
We build our approach upon two open-source LLMs, i.e., llama-2-7b666https://huggingface.co/meta-llama/Llama-2-7b-hf and bloomz-7b1-mt777https://huggingface.co/bigscience/bloomz-7b1-mt. We verify the effectiveness of our proposed approach on five translation tasks, including {Chinese (ZH), French (FR), German (DE), Spanish (ES), Russian (RU)}English (EN).
@@ 3.1 Experimental Settings
The corpus of all translation tasks is extracted from News-Commentary-v18. For LLM-based models, We use the tokenizer of foundation models to process the input data and no other preprocessing is performed. See Appendix A for more details on splitting, preprocessing and statistics of datasets. Besides, we provide a discussion for scales of the training set in Appendix I .
@@
In addition to traditional context-agnostic (Trans.) and context-aware (G-Trans Bao et al. (2021) and MR-Trans Sun et al. (2022) with or without pre-training setting, i.e., + mBART liu2020multilingual) NMT models with encoder-decoder architecture ,888Please refer to Appendix C for more introduction about the G-Trans and MR-Trans. our primary comparison focuses on the following three LLM-based alternatives: 1) MT-LoRA: It is a tuned LLM adapted to NMT task via the tuning method of Low-Rank Adaptation Hu et al. (2022), which makes large-scale pre-training models adapt to a new task by injecting a trainable rank decomposition matrice into each layer of the Transformer architecture; 2) MT-PT: It is a tuned LLM adapted to NMT task via the deep prompt tuning with MLPs reparameterization,999We attempt to remove reparameterization but experience a significant decline in performance. which only tunes continuous prompts with a frozen language model; 3) CMT-PT: It indiscriminately utilizes inter- and intra-sentence context via the concatenation strategy, as depicted in Figure 1 (a). Similar to MT-PT, it is also a tuned LLM via the deep prompt tuning with MLPs reparameterization. Among them, MT-LoRA and MT-PT are context-agnostic systems while CMT-PT is a context-aware system. For a fair comparison, we ensure that all context-aware models built upon LLM, including CMT-PT, MPT, and DeMPT, incorporate identical inter-sentence context. We provide more discussion in utilization of various inter-sentence contexts in Appendix L and M.
@@
For all encoder-decoder Transformer models, including Transformer (Trans.), MR-Trans and G-Trans101010For G-Trans, we use their official implementation upon Fairseq. Code: https://github.com/baoguangsheng/g-transformer., we implement them upon Fairseq Ott et al. (2019). For MT-LoRA models, we set the rank of trainable matrices as 16 which performs best in our preliminary experiment. For all MT-PT models, CMT-PT models, and our models, we set the prompt length as 64.111111We provide more discussion in Appendix E about the prompt length. For the incorporation of inter-sentence context in CMT-PT and our models, we consider a dynamic , in which the total tokens are no more than 256. In enhanced decoding, we consider the three next word predictions to be equally important by setting both and to 1/3. We provide an analysis of and more training details in Appendix K and B, respectively.
@@
We use sacreBLEU (accuracy-related metric)121212Signature: nrefs:1|case:mixed|eff:no|tok:13a|
smooth:exp|version:2.3.1 Post (2018), COMET (semantics-related metric) with the wmt22-comet-da model131313https://github.com/Unbabel/COMET Rei et al. (2020), and BlonDe (discourse-related metric) Jiang et al. (2022) as evaluation metrics.141414We provide more discourse-related evaluation in Appendix F.